# Data-Driven Integration Kernels for Interpretable Nonlocal Operator Learning

### 11 Mar 2026 v1

@@ Appendix B Neural Network Architecture and Training Configuration
All models in the hierarchy share an identical downstream neural network that maps either flattened predictor fields (baseline models) or kernel-integrated features (kernel-based models), together with local inputs, to precipitation. This ensures that differences in predictive performance across models reflect only differences in how nonlocal information is represented, not differences in network architecture or optimization.
The downstream network is a fully connected feedforward model with four hidden layers of widths 256, 128, 64, and 32. Each hidden layer is followed by a GELU activation and dropout with probability 0.1. The output layer consists of a single linear neuron and predicts precipitation in standardized, log-transformed space.
Models are trained with a batch size of 500 using the Adam optimizer and an initial learning rate of $5\times 10^{-4}$. The learning rate is reduced by a factor of 0.5 after two consecutive epochs without validation improvement, with a minimum learning rate of $10^{-6}$. Training proceeds for up to 20 epochs with early stopping based on validation loss, terminating after four consecutive non-improving epochs. The loss function is mean squared error in this standardized, log-transformed space, and all experiments use fixed random seeds (42, 72, 102) to ensure reproducibility.
@@ Appendix C Additional Results
Model $\mathbf{R^{2}\pm 1\sigma}$
$(\mathbf{x},\mathit{p},\mathit{t})$ $0.5765\pm 0.0017$
$(\mathbf{x}_{0},\mathit{p},\mathit{t}_{0})$ $0.5250\pm 0.0012$
$(\mathbf{x}_{0},\mathit{p}^{\mathit{k}},\mathit{t}_{0})$ $0.4970\pm 0.0002$
$(\mathbf{x}_{0},\mathit{p}^{\mathrm{MIX}},\mathit{t}_{0})$ $0.4885\pm 0.0010$
$(\mathbf{x}_{0},\mathit{p}^{\mathrm{MG}},\mathit{t}_{0})$ $0.4884\pm 0.0003$
$(\mathbf{x}_{0},\mathit{p}^{\mathrm{G}},\mathit{t}_{0})$ $0.4848\pm 0.0014$
$(\mathbf{x}_{0},\mathit{p}^{\mathrm{EXP}},\mathit{t}_{0})$ $0.4827\pm 0.0011$
$(\mathbf{x}_{0},\mathit{p}^{\mathrm{TH}},\mathit{t}_{0})$ $0.4820\pm 0.0013$
$(\mathbf{x},\mathit{p}_{0},\mathit{t}_{0})$ $0.4397\pm 0.0027$
$(\mathbf{x}_{0},\mathit{p}_{0},\mathit{t})$ $0.4334\pm 0.0004$
$(\mathbf{x}_{0},\mathit{p}_{0},\mathit{t}_{0})$ $0.4134\pm 0.0006$
Table A1: Mean test set $R^{2}\pm 1$ standard deviation across training seeds (n = 3) for all models shown in Figure 2. Values are computed in the standardized, log-transformed space
Figure A1: Learned vertical kernels for all kernel-based models shown in Figure 2, extending the results in Figure 3
@@ Acknowledgments
We gratefully acknowledge support from various funding sources. S.L.F. was supported by the National Science Foundation (NSF) Science and Technology Center (STC) Learning the Earth with Artificial Intelligence and Physics (LEAP) (2019625-STC) and the U.S. Department of Energy (DOE) Advanced Scientific Computing Research (ASCR) program (DE-SC0022255). J.L. was supported in part by the Program for Climate Model Diagnosis and Intercomparison (PCMDI), an Earth System Model Evaluation Project supported by the Earth and Environmental Systems Modeling (EESM) program in the DOE Office of Science (OS). S.S. was supported by the DOE Office of Biological and Environmental Research (BER) (DE-SC0025456). M.S.P. acknowledges support from NVIDIA Corporation. J.W.B. was supported by the National Aeronautics and Space Administration (NASA) New Investigator Program in Earth Science (NIP) (80NSSC21K1735). T.B. was supported by the Swiss State Secretariat for Education, Research, and Innovation (SERI) under the Horizon Europe AI4PEX Project (Grant Agreement ID 101137682; SERI No. 23.00546). Computational resources were provided by Perlmutter at the National Energy Research Scientific Computing Center (NERSC), a DOE User Facility (Project m4334).
@@
ERA5 [Hersbach2023a, Hersbach2023b] and IMERG V06 [Huffman2019] data used in this study are publicly available in analysis-ready, cloud-optimized formats via the LEAP Data Catalog (https://catalog.leap.columbia.edu/feedstock/arco-era5) and the Microsoft Planetary Computer Data Catalog (https://planetarycomputer.microsoft.com/dataset/gpm-imerg-hhr), respectively. All code used for data retrieval and processing, model training and inference, and visualization is publicly available at (https://github.com/savannah-ferretti/monsoon-kernels).
@@
T.B. developed the initial research concept and provided supervision throughout the project. S.L.F. conducted the investigation, including data curation, software development, formal analysis, validation, and visualization, while methodology was developed by T.B. with technical support from S.L.F. and J.L. Funding acquisition and project administration were led by T.B. and M.S.P., with additional administrative support from J.W.B. and resources provided by M.S.P. The manuscript was drafted by S.L.F. and T.B., and refined through critical review and editing by J.L., S.S., J.W.B., and M.S.P. All authors approved the final submitted draft.

### 31 Mar 2026 v2

@@ Appendix B Neural Network Architecture and Training Configuration
All models in the hierarchy share an identical downstream neural network that maps either flattened predictor fields (baseline models) or kernel-integrated features (kernel models), together with local inputs, to precipitation. This ensures that differences in predictive performance across models reflect only differences in how nonlocal information is represented, not differences in network architecture or optimization.
The downstream network is a fully connected feedforward model with four hidden layers of widths 256, 128, 64, and 32. Each hidden layer is followed by a GELU activation and dropout with probability 0.1. The output layer consists of a single linear neuron and predicts precipitation in standardized, log1p-transformed space.
Models are trained with a batch size of 500 using the Adam optimizer and an initial learning rate of $5\times 10^{-4}$. The learning rate is reduced by a factor of 0.5 after two consecutive epochs without validation improvement, with a minimum learning rate of $10^{-6}$. Training proceeds for up to 20 epochs with early stopping based on validation loss, terminating after four consecutive non-improving epochs. The loss function is MSE in this standardized, log1p-transformed space, and all experiments use fixed random seeds (42, 72, 102) to ensure reproducibility.
@@ Appendix C Additional Results
Model $\mathbf{R^{2}}$ $\mathbf{MSE}$
$(\mathbf{x}_{0},p_{0},t_{0})$ $0.411\pm 0.0015\,(0.0006)$ $0.680\pm 0.0039\,(0.0006)$
$(\mathbf{x}_{0},p_{0},t)$ $0.432\pm 0.0015\,(0.0002)$ $0.656\pm 0.0037\,(0.0003)$
$(\mathbf{x},p_{0},t_{0})$ $0.440\pm 0.0016\,(0.0028)$ $0.647\pm 0.0035\,(0.0033)$
$(\mathbf{x}_{0},p^{\mathrm{TH}},t_{0})$ $0.481\pm 0.0015\,(0.0014)$ $0.599\pm 0.0034\,(0.0016)$
$(\mathbf{x}_{0},p^{\mathrm{EXP}},t_{0})$ $0.482\pm 0.0015\,(0.0010)$ $0.598\pm 0.0035\,(0.0012)$
$(\mathbf{x}_{0},p^{\mathrm{G}},t_{0})$ $0.484\pm 0.0014\,(0.0013)$ $0.596\pm 0.0035\,(0.0016)$
$(\mathbf{x}_{0},p^{\mathrm{MG}},t_{0})$ $0.487\pm 0.0015\,(0.0003)$ $0.592\pm 0.0034\,(0.0004)$
$(\mathbf{x}_{0},p^{\mathrm{MIX}},t_{0})$ $0.488\pm 0.0014\,(0.0009)$ $0.592\pm 0.0033\,(0.0011)$
$(\mathbf{x}_{0},p^{k},t_{0})$ $0.496\pm 0.0015\,(0.0002)$ $0.582\pm 0.0033\,(0.0003)$
$(\mathbf{x}_{0},p,t_{0})$ $0.528\pm 0.0015\,(0.0013)$ $0.546\pm 0.0032\,(0.0015)$
$(\mathbf{x},p,t)$ $0.582\pm 0.0014\,(0.0017)$ $0.482\pm 0.0028\,(0.0019)$
Table A1: Test set performance of models in Figure 2, computed in the standardized log1p-transformed precipitation space. Results are reported as the mean across seeds, with bootstrap standard deviation ($n$ = 1,000 resamples) and across-seed standard deviation ($n$ = 3) in parentheses
Figure A1: Spatial composites of standardized, log1p-transformed precipitation by regime. Columns show the ground truth, baseline nonlocal vertical model, and our nonparametric and best parametric (MIX) kernel models. Rows correspond to dry ($\leq 5^{\mathrm{th}}$), typical (45–55th), and wet ($\geq 95^{\mathrm{th}}$) percentiles of domain-mean precipitation. Fields are averaged over selected timesteps within each regime
Figure A2: Learned vertical kernels for all kernel models, extending the results in Figure 3
@@ Acknowledgments
S.L.F. was supported by the National Science Foundation (NSF) Science and Technology Center (STC) Learning the Earth with Artificial Intelligence and Physics (LEAP) (2019625-STC) and the U.S. Department of Energy (DOE) Advanced Scientific Computing Research (ASCR) Program (DE-SC0022255). J.L. was supported by the DOE Office of Science (OS) through the Program for Climate Model Diagnosis and Intercomparison (PCMDI). S.S. was supported by the DOE Office of Biological and Environmental Research (BER) (DE-SC0025456). M.S.P. acknowledges support from NVIDIA Corporation. J.W.B. was supported by the National Aeronautics and Space Administration (NASA) New Investigator Program in Earth Science (NIP) (80NSSC21K1735). T.B. was supported by the Swiss State Secretariat for Education, Research, and Innovation (SERI) under the Horizon Europe AI4PEX Project (Grant Agreement ID 101137682; SERI No. 23.00546). Computational resources were provided by Perlmutter at the National Energy Research Scientific Computing Center (NERSC), a DOE User Facility (Project m4334).
@@
ERA5 [Hersbach2023a, Hersbach2023b] and IMERG V06 [Huffman2019] data used in this study are publicly available in analysis-ready, cloud-optimized formats via the LEAP Data Catalog (https://catalog.leap.columbia.edu/feedstock/arco-era5) and the Microsoft Planetary Computer Data Catalog (https://planetarycomputer.microsoft.com/dataset/gpm-imerg-hhr), respectively. All code used for data retrieval and processing, model training and inference, and visualization is available on GitHub (https://github.com/savannahferretti/monsoon-kernels).
@@
T.B. developed the initial research concept and provided supervision throughout the project. S.L.F. conducted the investigation, including data curation, software development, formal analysis, validation, and visualization, while the methodology was developed by T.B. with technical support from S.L.F. and J.L. Funding acquisition and project administration were led by T.B. and M.S.P., with additional administrative support from J.W.B. and resources provided by M.S.P. The manuscript was drafted by S.L.F. and T.B., and refined through critical review and editing by J.L., S.S., J.W.B., and M.S.P. All authors approved the final submitted draft.